# A Generic Trajectory Planning Method for Constrained All-Wheel-Steering Robots

### v2

@@ III-C Trajectory Optimization
where, $\mathbf{x}(t)$ represents the states at time $t$, $dt$ is a short time duration, $k_{1},k_{2},k_{3},k_{4}$ are estimated slopes by the ordinary differential equation $\frac{\partial\mathbf{x}}{\partial t}=\dot{\mathbf{x}}=\begin{bmatrix}&\dot{x}&\dot{y}&\dot{\theta}&\mathbf{u}^{\top}\end{bmatrix}^{\top}$, by this case, the continuous constrain between $\mathbf{x}(t)$ and $\mathbf{x}(t+dt)$ is
@@
Mode changing constraints: The control continuous space is divided into phases. The wheel rotation directions are provided by the initial guess. The rotating directions $D_{w}\in\{-1,1\}$ are searched according to the mapping table, minimum cost directions are selected if the steering ranges are overlapped. The determination of the direction of this round of rotation can be modeled as a Mixed Integer Nonlinear Optimization (MINLP) [31] problem, but its solution time is very long. The hybrid A* process reduced the online optimization burden that may be introduced by MINLP.
@@
In addition to this, due to the discontinuity in the speed state space when switching wheel modes. Hybrid A* also provides an indicator for manipulation phase $M\in\mathbb{N}^{H\times 1}$ transition keyframes. The first and last frames of different modes are set as keyframes, and the speed of this frame is 0, which can be expressed by the equation as
$G_{M}:=v_{t}*\left(|M(t)-M(t+dt)|+|M(t)-M(t-dt)|\right)=0.$
@@
where A is the diagonal coefficient matrix, $T$ is the total time, $\mathbf{u}(t)$ is the state vector, $\bar{\mathbf{x}}_{0}$ and $\bar{\mathbf{x}}_{f}$ are the initial and final states, H is the sampling steps in the time horizon, and $k_{t}$ is the weight of time optimization, F is an arbitrary task-related cost function. Here we make the distance between original waypoints and smoothed trajectory as the task-oriented cost function. Besides, wheel acceleration can be constrained with eq. 5 itself, avoiding complex computation in eq. 6.
@@ IV-A Constrained Trajectory Optimization
Refer to captionRefer to caption
Figure 6: The rear steering limits are set to ±$\pm 0.001$°and the front steering limits are both ±$\pm 75$°. The position of the rear wheel to the control center point is symmetrical with the front wheel in the left column. The position of the rear wheels is moved forward to both sides of the control center point, similar to the bicycle model in the right figure.
@@ IV-B Trajectory Following Experiments
We conducted the trajectory following experiments in NVIDIA Isaac Sim [34] simulator, which can offer a realistic physics engine and accurate robot status. The simulator allows us to thoroughly develop and test our trajectory generator with high repeatability and flexibility. These evaluations were primarily carried out in a car park environment as Fig. 7 where precise and agile maneuvers are essential. To minimize the impact of the controller, the path followers are the same as C-AWS trajectory generators except the time horizon is much shorter. Velocity commands of the vehicle body are subsequently translated into specific velocity and angular commands for each wheel in the Vehicle Control Unit (VCU) at a rate of $\sim 10Hz$.
@@
Figure 8: Smoothed paths and motion records are depicted in the simulator. The robots are tasked with locomoting from the pose on the right to the left. The black line represents the reference path, while the orange line indicates the recorded control center. Additionally, the orange and blue rectangles illustrate the recorded body and wheel poses, respectively.
@@
From the first row in Fig. 89, it is evident that due to the absence of foreseeing steering constraints in OMNI, the wheels experience substantial unexpected jittering and slippage at discontinuous points in the control space, which in turn negatively impacts both the accuracy of trajectory tracking and the execution speed. In the second row in Fig. 89, the optimal trajectory is generated using S-AWS, whose chassis geometry can also be considered as a standard bicycle model with a half-wheel base. From the figures, there are noticeable trade-offs as the vehicle’s posture adjustments consume a significant amount of time, thereby prolonging the overall execution time. Indeed, due to the matching rotation angles of both the front and rear wheels of this chassis, it is inherently less susceptible to slipping. By contrast, our C-AWS optimizer facilitates the robot with more flexibility and completes the route with less time, and lower tracking error, while also exhibiting reduced wheel slippage, which can be readily drawn from the third row in Fig. 89.
@@ IV-C Quantitative Comparison
TABLE I: Dynamic Performance Comparison(Units: m, s, rad)
Method $\Delta|d|$ $\Delta\theta$ slide ratio $\lambda$ ave. velocity ave. acceleration ave. jerk
@@
TABLE II: Operation Time Comparison(Units: s)
Method Success Rate Search 20steps Optimize 20steps Headless 20steps
@@
We quantitatively evaluated the aforementioned trajectory generator and followers, including OMNI: to prove the impact on vehicle motion if the steering constraint is ignored; S-AWS: to prove the generalization of our proposed method and that it can satisfy the basic model of the bicycle model; C-AWS: to prove it really works and illustrate its properties. We additionally add baseline w/o smooth by directly feeding the trajectory generated through hybrid A* to the trajectory tracking controller for tracking. The comparison with this baseline is mainly to prove the contribution of our proposed method to the smoothness of vehicle driving and the ease of trajectory tracking.
@@
From Table I we can see that the method we proposed has a lower wheel slide ratio, which can lead to more stable control and less mechanical damage to the vehicle during driving. And there is smaller acceleration during driving, which can better ensure the comfort of the passengers and the safety of the cargo carried. By comparing with w/o smooth, the transformation speed before optimization is faster, but this is achieved by sacrificing the acceleration and deceleration process and traceability performance. The operation times are listed in Table II. Compared with OMNI who despite the wheel constraints, our method illustrates that computation of wheel constraints will lead to a longer trajectory planning time, they can still basically meet the real-time requirements of the trajectory planning module. Furthermore, we computed the optimization time excluding the initialization time. The data reveals that our running time still retains approximately 40% optimization potential.
V Conclusion and Discussion
@@
In this work, we addressed the trajectory planning problem for constrained all-wheel steering (AWS) robots. We employed a search-first-then-optimize framework, reformulating the AWS kinematic model to tackle the nonlinearity arising from steering limits. In this framework, a time-cost-2nd-order initial trajectory planner was introduced. It was widely suitable for robots with fixed steering axis positions. Additionally, a trajectory optimizer that incorporates our kinematic model was proposed, linearizing steering constraints to generate feasible paths and control mechanisms for C-AWS. Our approach not only enhanced maneuverability but also improved passenger comfort by reducing acceleration changes. Compared to alternatives, our method was superior and met the requirements for real-time operation.
@@
For future work, we aim to enhance the efficiency of our method by optimizing the running speed of the search and optimization processes. We also plan to integrate our approach with various tasks to assess how this type of chassis can be improved for different applications. Additionally, the constrained AWS (C-AWS) system will be upgraded with modern differentiable collision-avoidance algorithms [29]. These improvements will be applicable in both static and dynamic environments, promising safer and more efficient transportation.

### v3

@@ III-C Trajectory Optimization
where $\mathbf{x}(t)$ represents the states at time $t$, $dt$ is a short time duration, $k_{1},k_{2},k_{3},k_{4}$ are estimated slopes by the ordinary differential equation $\frac{\partial\mathbf{x}}{\partial t}=\dot{\mathbf{x}}=\begin{bmatrix}&\dot{x}&\dot{y}&\dot{\theta}&\mathbf{u}^{\top}\end{bmatrix}^{\top}$, by this case, the continuous constrain between $\mathbf{x}(t)$ and $\mathbf{x}(t+dt)$ is
@@
Mode changing constraints: The control continuous space is divided into phases. The wheel rotation directions are provided by the initial guess. The rotating directions $D_{w}\in\{-1,1\}$ are searched according to the mapping table, minimum cost directions are selected if the steering ranges are overlapped. The determination of the direction of this round of rotation can be modeled as a Mixed Integer Nonlinear Programing (MINLP) [30] problem, but its solution time is very long. The Hybrid-A* process reduced the online optimization burden that may be introduced by MINLP.
@@
In addition to this, due to the discontinuity in the speed state space when switching wheel modes. Hybrid-A* also provides an indicator for manipulation phase $M\in\mathbb{N}^{H}$ transition keyframes. The first and last frames of different modes are set as keyframes, and the speed of this frame is 0, which can be expressed by the equation as
$G_{M}:=v_{t}\cdot\left(|M(t)-M(t+dt)|+|M(t)-M(t-dt)|\right)=0.$
@@
where $A$ is the diagonal coefficient matrix, $T$ is the total time, $\mathbf{u}(t)$ is the state vector, $\bar{\mathbf{x}}_{0}$ and $\bar{\mathbf{x}}_{f}$ are the initial and final states, $H$ is the sampling steps in the time horizon, and $k_{t}$ is the weight of time optimization, $F$ is an arbitrary task-related cost function. Here we define the $L2$ distance between poses of initial guess and smoothed trajectory as the task-oriented cost function. Besides, wheel acceleration can be constrained with Eq. (5) itself.
@@ IV-A Constrained Trajectory Optimization
Figure 6: The annotations in this figure is the same as Fig. 5. Specially, for both left and right columns, the rear steering limits are set to ±$\pm 0.001$° and the front steering limits are both ±$\pm 75$°. In the left column, the position of the rear wheel to the control center point is symmetrical with the front wheel. In the right column, the position of the rear wheels is moved forward besides the control center.
@@ IV-B Trajectory Following Experiments
[…]
@@
Figure 8: Smoothed paths and motion records are depicted in the simulator. The robots are tasked with locomotion from the pose on the right to the left. The black line represents the reference path, while the orange line indicates the recorded control center. Additionally, the orange and blue rectangles illustrate the recorded body and wheel poses, respectively.
@@
From the first row in Figs. 8 and 9, it is evident that due to the absence of foreseeing steering constraints in OMNI, the wheels experience substantial unexpected jittering and slippage at discontinuous points in the control space, which in turn negatively impacts both the accuracy of trajectory tracking and the execution speed. In the second row in Figs. 8 and 9, the optimal trajectory is generated using S-AWS, whose chassis geometry can also be considered as a standard bicycle model with a half-wheel base. From the figures, there are noticeable trade-offs as the vehicle’s posture adjustments consume a significant amount of time, thereby prolonging the overall execution time. Indeed, due to the matching rotation angles of both the front and rear wheels of this chassis, it is inherently less susceptible to slipping. By contrast, our C-AWS optimizer facilitates the robot with more flexibility and completes the route with less time, and lower tracking error, while also exhibiting reduced wheel slippage, which can be readily drawn from the third row in Figs. 8 and 9.
@@ IV-C Quantitative Comparison
TABLE I: Dynamic Performance Comparison
[…]
@@
TABLE II: Operation Time Comparison
Method Success Rate Search 20steps (s) Optimize 20steps (s) Headless 20steps (s)
@@
We quantitatively evaluated the aforementioned trajectory generator and followers, including OMNI: to prove the impact on vehicle motion if the steering constraint is ignored; S-AWS: to prove the generalization of our proposed method and that it can satisfy the basic model of the bicycle model; C-AWS: to prove it really works and illustrate its properties. We additionally add baseline W/O Smooth by directly feeding the trajectory generated through Hybrid-A* to the trajectory tracking controller for tracking. The comparison with this baseline is mainly to prove the contribution of our proposed method to the smoothness of vehicle driving and the ease of trajectory tracking. Average tracking error $\Delta|d|$, average heading error $\Delta\theta$, and the maximum slide ratio $\lambda$ among wheels and dynamic performance are listed in Table I.
@@
From Table I, we can see that the method we proposed has a lower wheel slide ratio, which can lead to more stable control and less mechanical damage to the vehicle during driving. And there is smaller acceleration during driving, which can better ensure the comfort of the passengers and the safety of the cargo carried. By comparing with W/O Smooth, the transformation speed before optimization is faster, but this is achieved by sacrificing the acceleration and deceleration process and traceability performance. The operation times are listed in Table II. Our method illustrates that computation of wheel constraints will lead to a longer trajectory planning time, they can still basically meet the real-time requirements of the trajectory planning module. Furthermore, we computed the optimization time excluding the initialization time of optimizer which can be accomplished by programming refraction. The results are labeled as Headless and reveals there still retains approximately 40% operation time reduction potential.
V Conclusion
@@
In this work, we addressed the trajectory planning problem for C-AWS robots. We employed a search-first-then-optimize framework, reformulating the AWS kinematic model to tackle the nonlinearity arising from steering limits. In this framework, a time-cost-2nd-order initial trajectory planner was introduced. It was widely suitable for robots with fixed steering axis positions. Additionally, a trajectory optimizer that incorporates our kinematic model was proposed, linearizing steering constraints to generate feasible paths and control mechanisms for C-AWS. Our approach not only enhanced maneuverability but also improved passenger comfort by reducing acceleration changes. Compared to alternatives, our method was superior and met the requirements for real-time operation.
@@
For future work, we aim to enhance the efficiency of our method by optimizing the running speed of the search and optimization processes. We also plan to integrate our approach with various tasks to assess how this type of chassis can be improved for different applications. Additionally, the C-AWS system will be upgraded with modern differentiable collision-avoidance algorithms [28]. These improvements will be applicable in both static and dynamic environments, promising safer and more efficient transportation.